OpenStudy (anonymous):

sqrt (72)- sqrt(8)+sqrt(128)

OpenStudy (anonymous):

You need to get everything with the same base radical in order to add them.

OpenStudy (anonymous):

hint: start with simplifying the middle term first

OpenStudy (anonymous):

thats 2 right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sqrt(72) = sqrt (36 * 2) = 6 sqrt(2) sqrt(8) = sqrt(4 * 2) = 2 sqrt(2) sqrt(128) = sqrt(64 * 2) = 8 sqrt(2) 6sqrt(2) + 2sqrt(2) + 8sqrt(2) = 16 sqrt(2) When adding radicals, you have to find a way to make all of the radicals the same. This doesn't always happen!

OpenStudy (anonymous):

the cube root of 8 is 2. The square root of 8 is \[2\sqrt{2}\]

OpenStudy (anonymous):

well yes that

OpenStudy (anonymous):

cause 2*2*2=8

OpenStudy (anonymous):

@smehta9 so the answer would be 12 sqrt(2) right

OpenStudy (anonymous):

oh yes, you're right. I'm sorry, I saw the minus sign as a plus sign.

OpenStudy (anonymous):

o ok thank you lol